# Technische Universität München, D-85748 Garching, Germany

### v1

@@ 1 Introduction
Theoretical predictions for $b\bar{b}H$ production at the LHC rely on two main approaches for treating the bottom-quark mass: the five-flavor scheme (5FS) and the four-flavor scheme (4FS). In the 5FS, the bottom quark is treated as a massless parton, with logarithmic contributions of collinear origin resummed into the parton distribution functions (PDFs). This assumption simplifies calculations in the 5FS, c.f. the leading-order (LO) diagram in figure 1 (a), allowing higher-order corrections in the strong coupling constant to be more readily computed. Substantial progress has been made in recent years on contributions to the cross section proportional to $y_{b}^{2}{}$ in the 5FS Das:2023rif ; Dicus:1998hs ; Balazs:1998sb ; Harlander:2003ai ; Campbell:2002zm ; Harlander:2010cz ; Ozeren:2010qp ; Harlander:2011fx ; Buehler:2012cu ; Belyaev:2005bs ; Harlander:2014hya ; Ahmed:2014pka ; Gehrmann:2014vha ; Duhr:2019kwi ; Mondini:2021nck ; Wiesemann:2014ioa ; Krauss:2016orf ; Ajjath:2019ixh ; Ajjath:2019neu ; Forte:2019hjc ; Badger:2021ega ; Das:2024pac . Notably, the computation of the third-order QCD cross section Duhr:2019kwi represents a significant milestone. Pure QED and mixed QCD-QED corrections in this framework are minimal, typically below $0.03\%$ of the LO cross section AH:2019xds , while mixed QCD-electroweak corrections contribute around $1\%$ Pagani:2020rsg . More recently, the matching of NNLO QCD with parton showers (NNLO+PS) has been performed in the 5FS in ref. Biello:2024vdh by some of us.
@@ 2 Outline of the calculation
where the final state is inclusive over the radiation of additional particles $X$. Contributions from the loop-induced $ggF$ process proportional to $y_{t}$, as illustrated in figure 1 (c), are excluded throughout this paper. Instead, our focus is on the terms proportional to $y_{b}^{2}$, which form a gauge-invariant subset of the cross section, which we shall refer to as $b\bar{b}H$ production for the remainder of this manuscript. In our calculation of the $b\bar{b}H$ process we consider the bottom quarks to be massive, i.e. we employ the 4FS, see figure 1 (b) for a representative LO diagram.
@@ 3.1 Original MiNNLOPS method
In the following, we summarise the MiNNLOPS method for colour-singlet production, which was initially introduced in refs. Monni:2019whf ; Monni:2020nks and has been applied to several processes by now Lombardi:2020wju ; Lombardi:2021rvg ; Buonocore:2021fnj ; Lombardi:2021wug ; Zanoli:2021iyp ; Gavardi:2022ixt ; Haisch:2022nwz ; Lindert:2022qdd ; Biello:2024vdh . In the following, we briefly review the method, and we refer to refs. Monni:2019whf ; Monni:2020nks for further details.
@@ 3.2 Extension to ${Q\bar{Q}}$ processes
The MiNNLOPS approach is currently the only NNLO+PS method that also extends to processes involving colour charges in initial and final state, including heavy-quark pair ($Q\bar{Q}$) production Mazzitelli:2020jio ; Mazzitelli:2021mmm ; Mazzitelli:2023znt and, very recently, the production of a heavy quark pair in association with colour-singlet particles ($Q\bar{Q}$) Mazzitelli:2024ura . In the following, we briefly recall these extensions of the MiNNLOPS method, which exploits the knowledge of the singular structure of the cross section at small transverse momentum of the final-state system. The structure of large logarithmic contributions for the $Q\bar{Q}$ final state Catani:2021cbl is very similar to the one valid for $Q\bar{Q}$ production at small $p_{\text{T}}{}$ Zhu:2012ts ; Li:2013mia ; Catani:2014qha ; Catani:2018mei . However, the $Q\bar{Q}$ case involves more general kinematic configurations compared to the $Q\bar{Q}$ case, where the heavy quarks are constrained to be back-to-back to each other in the Born configuration. In either case, the starting point is the following factorisation theorem, which is expressed in the Fourier-conjugate space to $p_{\text{T}}$ (so-called impact-parameter or $b$-space) Zhu:2012ts ; Li:2013mia ; Catani:2014qha ; Catani:2018mei :
@@ 4.1 Massification procedure
where $\mathcal{J}$, $\mathcal{S}$, and $\mathcal{H}$ are, respectively, the jet, soft and hard functions for massive bottom quarks. We stress that $\mathcal{M}^{\textrm{IR-div}}_{{c\bar{c}}}$ is the amplitude before the subtraction of the IR divergences in the small-mass limit, performed in the minimal way via the operator $\mathbf{Z}_{m_{b}\ll\mu_{Q},{c\bar{c}}}$ in order to obtain the finite remainder $|\mathcal{R}_{c\bar{c}}\rangle=\mathbf{Z}^{-1}_{m_{b}\ll\mu_{Q},{c\bar{c}}}|\mathcal{M}_{c\bar{c}}\rangle$. There is a freedom in organising sub-leading soft terms into the jet and soft functions. The following requirement can completely fix this ambiguity,
@@ 4.2 Numerical implementation of the massless two-loop amplitude
We have implemented a C++ library for a fast numerical evaluation of two-loop virtual corrections in the leading-colour approximation with massless bottom quarks.  [^2] In the very final stages of this work, the full-colour amplitude with massless bottom quarks was presented in ref. Badger:2024awe . The implementation of subleading contributions in our generator is left for future investigations. We have used the analytic results of ref. Badger:2021ega , where the authors provided the finite remainder $\mathcal{F}^{(2)}$,
@@
We have conducted cross-checks with an independent implementation described in ref. Devoto:2024nhl , using the Yukawa coupling renormalised in the on-shell scheme. After pointwise validation of the massless amplitudes in eq. (51) at selected random phase-space points for LHC energies in both the gluon and quark channels, we have successfully compared the two-loop finite remainder in the small-mass limit with the generalised approach outlined in eq. (44).
@@ 5.2 Inclusive cross section
4FS MiNNLOPS ($H_{T}/4$) $0.496(6)_{-14\%}^{+16\%}$ 1.289
5FS NLO+PS ($m_{H}$) $0.676(6)_{-15\%}^{+27\%}$ 1.000
5FS MiNLO ($m_{H}$) $0.571(1)_{-23\%}^{+17\%}$ 0.844
5FS MiNNLOPS ($m_{H}$) $0.509(1)_{-5.3\%}^{+2.9\%}$ 0.752
@@
We begin the phenomenological analysis of the $b\bar{b}H$ process by studying the total inclusive cross section in table 1 at NLO+PS (Powheg) and NNLO+PS (MiNNLOPS) in the 4FS and the 5FS, where we consider two scale choices for the bottom-Yukawa coupling, ${\mu_{\text{R}}^{(0),y}}=m_{H}$ and ${\mu_{\text{R}}^{(0),y}}=H_{T}/4$ in the 4FS. To facilitate this analysis, we have developed a private NLO+PS generator in Powheg-Box-Res for the $y_{b}^{2}$ contribution in the 4FS. [^3] To ensure consistency, the NLO+PS implementation has been cross-checked against the public version in Powheg-Box-V2 Jager:2015hka . The NLO+PS predictions are obtained using the same setup of the MiNNLOPS as discussed earlier, except for the scale variation of the Yukawa coupling, which is performed using the two-loop running. The factorisation and renormalisation central scales are set to the invariant mass of the $b\bar{b}H$ system in the NLO+PS generator. Using the setup of MiNNLOPS, we also provide MiNLO predictions, which are formally NLO accurate for this observable, by turning off the term $(D-D_{1}-D_{2})$ in eq. (3.1).
@@
Based on the MiNNLOPS predictions in table 1, the NNLO corrections increase the NLO cross section by $30\%$ for both Yukawa scales, making them essential for achieving precise predictions in the 4FS. Furthermore, the MiNNLOPS prediction has a relatively small sensitivity to the considered scales of the bottom-Yukawa coupling, especially considering that its square is an overall factor to the cross section, highlighting a reduced dependence on scale choice at NNLO. For comparison, we also show 5FS predictions (i.e. the process $b\bar{b}\rightarrow H+X$) in table 1.444The NLO+PS 5FS results presented in table 1 are obtained using the same setup as detailed in ref. Biello:2024vdh , incorporating four-loop running to obtain the central bottom-Yukawa coupling and three-loop running for scale variations. Similarly, the MiNNLOPS predictions in the 5FS are generated using the identical setup as described in ref. Biello:2024vdh . The cross section in 5FS at NNLO+PS appears to underestimate theoretical uncertainty due to scale variation, whereas the 4FS prediction provides a more conservative estimate. The MiNNLOPS results clearly demonstrate agreement between the two schemes within scale uncertainties. This highlights that the long-standing discrepancy between 4FS and 5FS predictions is resolved by the newly-computed NNLO corrections in the 4FS. A detailed comparison between the massive and massless schemes at differential level is presented in section 6.
@@ 6.2 Integrated results
$pp\rightarrow H+\geq 1\,b$ jets 5FS NLO+PS $92.(7)_{-15\%}^{+27\%}$ 1.000 1.000
5FS MiNNLOPS $104.(9)_{-8.7\%}^{+8.8\%}$ 1.132 1.000
4FS NLO+PS $69.(4)_{-16\%}^{+21\%}$ 1.000 0.749
4FS MiNNLOPS $92.(0)_{-12\%}^{+9.8\%}$ 1.326 0.877
$pp\rightarrow H+\geq 2\,b$ jets 5FS NLO+PS $4.8(9)_{-15\%}^{+27\%}$ 1.000 1.000
5FS MiNNLOPS $7.6(4)_{-10\%}^{+11\%}$ 1.562 1.000
4FS NLO+PS $4.4(9)_{-18\%}^{+25\%}$ 1.000 0.991
4FS MiNNLOPS $6.3(9)_{-9.1\%}^{+1.5\%}$ 1.423 0.836
$pp\rightarrow H+\geq 1\,\ell$ jets 5FS NLO+PS $32.(8)_{-15\%}^{+27\%}$ 1.000 1.000
5FS MiNNLOPS $53.(0)_{-7.6\%}^{+6.9\%}$ 1.616 1.000
4FS NLO+PS $31.(4)_{-18\%}^{+26\%}$ 1.000 0.987
4FS MiNNLOPS $38.(3)_{-12\%}^{+1.9\%}$ 1.220 0.722
@@
Interestingly, we notice that for the $\geq$2-$b$-jet and $-\ell$-jet the NLO+PS results in 4FS and 5FS are very close, even closer than for the MiNNLOPS predictions. These observables become successively less accurate in the 5FS, which at Born level does not feature bottom quarks in the hard matrix elements. Therefore, we conclude that the agreement at NLO+PS for these observables that are less accurate in the 5FS is completely accidental. Indeed, we will see when considering differential distributions in the next subsection that the shapes in 4FS and 5FS are vastly different at NLO+PS, which confirms the accidental agreement at the integrated cross-section level.
@@ 6.3 Differential results
Figure 9: Comparison of the two MiNNLOPS generators in the 5FS and 4FS with analytic results in the 5FS. The Higgs $p_{\text{T}}$ spectrum (left) is compared to the NNLO+NNLL predictions of ref. Harlander:2014hya using NNLO NNPDF 4.0 sets, while the Higgs rapidity spectrum (right) is compared with the NNLO fixed-order result of ref. Mondini:2021nck based on NNLO CT14 sets.
@@
In the right plot of figure 9, the Higgs rapidity distribution from the two MiNNLOPS generators are in full agreement with the NNLO fixed-order results from ref. Mondini:2021nck , which employs the NNLO set of the CT14 PDFs Dulat:2015mca . The agreement is particularly good in the central rapidity region.
@@ 8 Conclusions
With the completion of the two NNLO+PS generators in the four- and five-flavour schemes, it will be interesting to consider their consistent combination, since each scheme describes different contributions in bottom-quark mass more accurately. Thus, developing a unified approach that combines the two schemes at NNLO+PS level, which has never been done before, could achieve unprecedented precision in the description of associated $b\bar{b}H$ production and also provide a pathway to apply such approach to similar processes in the future. Additionally, our calculation provides the basis to develop MiNNLOPS generators for processes with lighter quark flavours in the final state, such as Higgs production in association with charm quarks in the final state. Not least, the gluon-fusion component to $b\bar{b}H$ production proportional to the top-Yukawa coupling could be considered with similar techniques. All these developments are left for future work.

### v2

@@ 1 Introduction
Theoretical predictions for $b\bar{b}H$ production at the LHC rely on two main approaches for treating the bottom-quark mass: the five-flavor scheme (5FS) and the four-flavor scheme (4FS). In the 5FS, the bottom quark is treated as a massless parton, with logarithmic contributions of collinear origin resummed into the parton distribution functions (PDFs). This assumption simplifies calculations in the 5FS, c.f. the leading-order (LO) diagram in figure 1 (a), allowing higher-order corrections in the strong coupling constant to be more readily computed. Substantial progress has been made in recent years on contributions to the cross section proportional to $y_{b}^{2}{}$ in the 5FS Das:2023rif ; Dicus:1998hs ; Balazs:1998sb ; Harlander:2003ai ; Campbell:2002zm ; Harlander:2010cz ; Ozeren:2010qp ; Harlander:2011fx ; Buehler:2012cu ; Belyaev:2005bs ; Harlander:2014hya ; Ahmed:2014pka ; Gehrmann:2014vha ; Duhr:2019kwi ; Mondini:2021nck ; Wiesemann:2014ioa ; Krauss:2016orf ; Ajjath:2019ixh ; Ajjath:2019neu ; Forte:2019hjc ; Badger:2021ega ; Das:2024pac ; Cal:2023mib . Notably, the computation of the third-order QCD cross section Duhr:2019kwi represents a significant milestone. Pure QED and mixed QCD-QED corrections in this framework are minimal, typically below $0.03\%$ of the LO cross section AH:2019xds , while mixed QCD-electroweak corrections contribute around $1\%$ Pagani:2020rsg . More recently, the matching of NNLO QCD with parton showers (NNLO+PS) has been performed in the 5FS in ref. Biello:2024vdh by some of us.
@@ 2 Outline of the calculation
where the final state is inclusive over the radiation of additional particles $X$. Contributions from the loop-induced $ggF$ process proportional to $y_{t}$, as illustrated in figure 1 (c), are excluded throughout this paper. Instead, our focus is on the terms proportional to $y_{b}^{2}$, which form a gauge-invariant subset of the cross section, which we shall refer to as $b\bar{b}H$ production for the remainder of this manuscript. For convenience, we leave the inclusion of the $y_{t}$-dependent contributions to future work, as these terms can be treated separately in perturbation theory. However, the leading-order $y_{t}^{2}$ contributions are already accounted for in NNLO calculations of the $ggF$ process. In our calculation of the $b\bar{b}H$ process we consider the bottom quarks to be massive, i.e. we employ the 4FS, see figure 1 (b) for a representative LO diagram.
@@ 3.1 Original MiNNLOPS method
In the following, we summarise the MiNNLOPS method for colour-singlet production, which was initially introduced in refs. Monni:2019whf ; Monni:2020nks and has been applied to several processes by now Lombardi:2020wju ; Lombardi:2021rvg ; Buonocore:2021fnj ; Lombardi:2021wug ; Zanoli:2021iyp ; Gavardi:2022ixt ; Haisch:2022nwz ; Lindert:2022qdd ; Biello:2024vdh . In the following, we briefly review the method, and we refer to refs. Monni:2019whf ; Monni:2020nks ; Ebert:2024zdj for further details.
@@ 3.2 Extension to ${Q\bar{Q}}$ processes
The MiNNLOPS approach is currently the only NNLO+PS method that also extends to processes involving colour charges in initial and final state, including heavy-quark pair ($Q\bar{Q}$) production Mazzitelli:2020jio ; Mazzitelli:2021mmm ; Mazzitelli:2023znt and, very recently, the production of a heavy quark pair in association with colour-singlet particles ($Q\bar{Q}$) Mazzitelli:2024ura . [^1] First steps towards the extension to processes with light jets in the final state has been made in ref. Ebert:2024zdj . In the following, we briefly recall these extensions of the MiNNLOPS method, which exploits the knowledge of the singular structure of the cross section at small transverse momentum of the final-state system. The structure of large logarithmic contributions for the $Q\bar{Q}$ final state Catani:2021cbl is very similar to the one valid for $Q\bar{Q}$ production at small $p_{\text{T}}{}$ Zhu:2012ts ; Li:2013mia ; Catani:2014qha ; Catani:2018mei . However, the $Q\bar{Q}$ case involves more general kinematic configurations compared to the $Q\bar{Q}$ case, where the heavy quarks are constrained to be back-to-back to each other in the Born configuration. In either case, the starting point is the following factorisation theorem, which is expressed in the Fourier-conjugate space to $p_{\text{T}}$ (so-called impact-parameter or $b$-space) Zhu:2012ts ; Li:2013mia ; Catani:2014qha ; Catani:2018mei :
@@ 4.1 Massification procedure
where $\mathcal{J}$, $\mathcal{S}$, and $\mathcal{H}$ are, respectively, the jet, soft and hard functions for massive bottom quarks. We stress that $\mathcal{M}^{\textrm{IR-div}}_{{c\bar{c}}}$ is the amplitude before the subtraction of the IR divergences in the small-mass limit, performed in the minimal way via the operator $\mathbf{Z}_{m_{b}\ll\mu_{Q},{c\bar{c}}}$ in order to obtain the finite remainder $|\mathcal{R}_{c\bar{c}}\rangle=\mathbf{Z}^{-1}_{m_{b}\ll\mu_{Q},{c\bar{c}}}|\mathcal{M}_{c\bar{c}}\rangle$. There is a freedom in organising subleading soft terms into the jet and soft functions. The following requirement can completely fix this ambiguity,
@@ 4.2 Numerical implementation of the massless two-loop amplitude
We have implemented a C++ library for a fast numerical evaluation of two-loop virtual corrections in the leading-colour approximation with massless bottom quarks.  [^3] In the very final stages of this work, the full-colour amplitude with massless bottom quarks was presented in ref. Badger:2024awe . The implementation of subleading contributions in our generator is disccused in Appendix B. We have used the analytic results of ref. Badger:2021ega , where the authors provided the finite remainder $\mathcal{F}^{(2)}$,
@@
We have conducted cross-checks with an independent implementation described in ref. Devoto:2024nhl , using the Yukawa coupling renormalised in the on-shell scheme. After pointwise validation of the massless amplitudes in eq. (51) at selected random phase-space points for LHC energies in both the gluon and quark channels, we have successfully compared the two-loop finite remainder in the small-mass limit with the generalised approach outlined in eq. (44). Additional cross-checks have been performed using the public library described in ref. Badger:2024awe , where we compared the numerical results of the full-colour amplitude in the large $N_{C}$ limit with those obtained from our leading-colour implementation.
@@ 5.2 Inclusive cross section
4FS MiNNLOPS ($H_{T}/4$) $0.494(6)_{-14\%}^{+16\%}$ 1.284
5FS NLO+PS ($m_{H}$) $0.645(5)_{-10\%}^{+11\%}$ 1.000
5FS MiNLO ($m_{H}$) $0.571(1)_{-23\%}^{+17\%}$ 0.885
5FS MiNNLOPS ($m_{H}$) $0.509(1)_{-5.3\%}^{+2.9\%}$ 0.790
@@
We begin the phenomenological analysis of the $b\bar{b}H$ process by studying the total inclusive cross section in table 1 at NLO+PS (Powheg) and NNLO+PS (MiNNLOPS) in the 4FS and the 5FS, where we consider two scale choices for the bottom-Yukawa coupling, ${\mu_{\text{R}}^{(0),y}}=m_{H}$ and ${\mu_{\text{R}}^{(0),y}}=H_{T}/4$ in the 4FS. To facilitate this analysis, we have developed a private NLO+PS generator in Powheg-Box-Res for the $y_{b}^{2}$ contribution in the 4FS. [^4] To ensure consistency, the NLO+PS implementation has been cross-checked against the public version in Powheg-Box-V2 Jager:2015hka . The NLO+PS predictions are obtained using the same setup of the MiNNLOPS as discussed earlier, except for the scale variation of the Yukawa coupling, which is performed using the two-loop running. The factorisation and renormalisation central scales are set to $H_{T}/4$ in the NLO+PS generator. Using the setup of MiNNLOPS, we also provide MiNLO predictions, which are formally NLO accurate for this observable, by turning off the term $(D-D_{1}-D_{2})$ in eq. (3.1).
@@
Based on the MiNNLOPS predictions in table 1, the NNLO corrections increase the NLO cross section by $30\%$ for both Yukawa scales, making them essential for achieving precise predictions in the 4FS. Furthermore, the MiNNLOPS prediction has a relatively small sensitivity to the considered scales of the bottom-Yukawa coupling, especially considering that its square is an overall factor to the cross section, highlighting a reduced dependence on scale choice at NNLO. For comparison, we also show 5FS predictions (i.e. the process $b\bar{b}\rightarrow H+X$) in table 1.555The NLO+PS 5FS results presented in table 1 are obtained using the same setup as detailed in ref. Biello:2024vdh , incorporating four-loop running to obtain the central bottom-Yukawa coupling and two-loop running for scale variations. Similarly, the MiNNLOPS predictions in the 5FS are generated using the identical setup as described in ref. Biello:2024vdh . The cross section in 5FS at NNLO+PS exhibits a smaller theoretical uncertainty from scale variation compared to the 4FS prediction, which provides a more conservative estimate. Different sources contribute to the distinct scale uncertainties of the two predictions, such as the presence of Born-level strong couplings and the sizable NNLO corrections in the massive calculation. The MiNNLOPS results clearly demonstrate agreement between the two schemes within scale uncertainties. This highlights that the long-standing discrepancy between 4FS and 5FS predictions is resolved by the newly-computed NNLO corrections in the 4FS. A detailed comparison between the massive and massless schemes at differential level is presented in section 6.
@@ 6.2 Integrated results
$pp\rightarrow H+\geq 1\,b$ jets 5FS NLO+PS $88.(1)_{-10\%}^{+11\%}$ 1.000 1.000
5FS MiNNLOPS $104.(9)_{-8.7\%}^{+8.8\%}$ 1.182 1.000
4FS NLO+PS $69.(4)_{-16\%}^{+21\%}$ 1.000 0.784
4FS MiNNLOPS $92.(1)_{-12\%}^{+9.8\%}$ 1.327 0.878
$pp\rightarrow H+\geq 2\,b$ jets 5FS NLO+PS $4.7(8)_{-11\%}^{+11\%}$ 1.000 1.000
5FS MiNNLOPS $7.6(4)_{-10\%}^{+11\%}$ 1.617 1.000
4FS NLO+PS $4.4(9)_{-18\%}^{+25\%}$ 1.000 0.936
4FS MiNNLOPS $6.4(4)_{-9.1\%}^{+1.5\%}$ 1.434 0.843
$pp\rightarrow H+\geq 1\,\ell$ jets 5FS NLO+PS $31.(1)_{-9.7\%}^{+12\%}$ 1.000 1.000
5FS MiNNLOPS $53.(0)_{-7.6\%}^{+6.9\%}$ 1.710 1.000
4FS NLO+PS $31.(4)_{-18\%}^{+26\%}$ 1.000 1.000
4FS MiNNLOPS $38.(3)_{-12\%}^{+2.0\%}$ 1.220 0.722
@@
Interestingly, we notice that for the $\geq$2-$b$-jet and $-\ell$-jet the NLO+PS results in 4FS and 5FS are very close, even closer than for the MiNNLOPS predictions. These observables become successively less accurate in the 5FS, which at Born level does not feature bottom quarks in the hard matrix elements. Therefore, we conclude that the agreement at NLO+PS for these observables that are less accurate in the 5FS is completely accidental. Indeed, we will see when considering differential distributions in the next subsection that the shapes in 4FS and 5FS are vastly different at NLO+PS, which confirms the accidental agreement at the integrated cross section level.
@@ 6.3 Differential results
Figure 9: Comparison of the two MiNNLOPS generators in the 5FS and 4FS with analytic results in the 5FS. The Higgs $p_{\text{T}}$ spectrum (left) is compared to the NNLO+NNLL predictions of ref. Harlander:2014hya using NNLO NNPDF 4.0 sets, while the Higgs rapidity spectrum (right) is compared with the NNLO fixed-order result of ref. Mondini:2021nck based on NNLO CT14 sets. For the right plot, we used NNLO CT14 sets in our MiNNLOPS generators, both in the 5FS and 4FS.
@@
In the right plot of figure 9, the Higgs rapidity distribution from the two MiNNLOPS generators are in full agreement with the NNLO fixed-order results from ref. Mondini:2021nck , which employs the NNLO set of the CT14 PDFs Dulat:2015mca  [^8] We thank the authors of ref. Mondini:2021nck for providing us with the relevant results.. Note that for this particular comparison, we used the CT14 PDF sets in our MiNNLOPS generators, both in the 5FS and 4FS. The agreement is particularly good in the central rapidity region. From this point onward, we revert to the default setting, using the NNPDF4.0 set for all subsequent simulations.
@@ 8 Conclusions
With the completion of the two NNLO+PS generators in the four- and five-flavour schemes, it will be interesting to consider their consistent combination, since each scheme describes different contributions in bottom-quark mass more accurately. The 5FS and 4FS MiNNLOPS generators presented in this work can indeed be used to obtain combined 4FS-5FS results. In order to perform the combination, we need to address the overlapping contributions between these two schemes, which requires modifying the current codes to ensure a consistent matching of the 4FS and 5FS. Thus, developing a unified approach, along the lines of refs. Gauld:2021zmq ; Guzzi:2024can for instance, that combines the two schemes at NNLO+PS level, which has never been done before, could achieve unprecedented precision in the description of associated $b\bar{b}H$ production and also provide a pathway to apply such approach to similar processes in the future. Additionally, our calculation provides the basis to develop MiNNLOPS generators for processes with lighter quark flavours in the final state, such as Higgs production in association with charm quarks in the final state. Not least, the gluon-fusion component to $b\bar{b}H$ production proportional to the top-Yukawa coupling could be considered with similar techniques. All these developments are left for future work.